Solve the Trigonometric equation in the interval [0, 2pi). Give the exact value, if possible; otherwise round your answer to two decimal places.

2cos^2θ + 9cosθ +4 = 0
Answer #1

2 cos?o +9 coso +4=0 2 cos 0 + 8 cos 0+ cos 0 +4=0 - 2coso (cos0+4) +1(cos 0+1) =0 > 2cos 0+1) (Cos 0+4) 0 = 2 cost=0 or coso